### v1

@@ 3.2 CTA and ASTRI IRFs
Fig. 1 also shows the IACT IRFs for on-axis observations (while IRFs for off-axis observations at different offsets from the FoV center are shown in Appendix A) for both ASTRI and the two CTA observatories. The latter can be retrieved in publicly available zenodo repositories (ASTRI-Project, 2022; Cherenkov Telescope Array Observatory & Cherenkov Telescope Array Consortium, 2021). Nominal pointing is assumed in the simulations, with all telescopes pointing to the same direction; different configurations are currently under investigation by the Collaborations.
A direct comparison among LHAASO, ASTRI, and CTA IRFs is now possible given that the same energy binning was set, namely 0.2 in logarithmic scale. The excellent angular resolution of IACTs, due to their pointing technique and camera, is visibly superior to the one of LHAASO-KM2A: in fact, the width of Point Spread Function (PSF) amounts to for IACTs vs for LHAASO. With regards to the effective area, instead, for EAS arrays, it coincides with the geometrical area, namely the instrument surface exposed to the shower, since shower reconstruction relies on the number of particles collected on ground. In the case of LHAASO, a surface of 1 km is covered by the KM2A. On the other hand, the effective area of IACTs is rather a convolution between the geometrical area and the Cherenkov light pool illuminated by the gamma-ray induced shower. The former depends on the specific layout of the experiment, namely by the relative distance among telescopes. For ASTRI, the foreseen layout covers an area of about m (Vercellone et al., 2022a), while CTA-North of about m and CTA-South of about m (Cherenkov Telescope Array Consortium et al., 2019). On the other hand, the Cherenkov pool of a primary photon interacting at a height of about 10 km spreads over a surface on ground of m (Aharonian et al., 2013): this value hence represents the effective area of an individual imaging telescope, which has to be rescaled to the number of telescopes when providing a description of the array. The effective areas shown in Fig. 1 are obtained after the gamma/hadron separation cuts, while no selection relative to the reconstructed event direction was applied333https://doi.org/10.5281/zenodo.5499840.
@@ 4 Sensitivity studies towards extended sources
The careful evaluation of the instrument sensitivities is important for obtaining a clear understanding of the expected performance and discovery potential, as discussed for ground-based telescopes by Aharonian (1991). We here calculate differential sensitivities for extended sources of different instruments following the approach described in Ambrogi et al. (2018). Namely, we identify the minimum detectable flux that guarantees the following conditions to be satisfied in each energy bin of the computation (Bernlöhr et al., 2013), namely:
@@
The number of signal events expected per energy bin is obtained by folding the gamma-ray effective area with the Crab nebula spectrum444This is described in the TeV regime as a power-law function, , where (TeV cm s) is the flux normalization at 1 TeV and (Aharonian et al., 2004).. However, it should be noted that the final sensitivity computation is independent of the assumed source flux. The number of background events is computed from the background rate, assumed to be constant in time. While the background rate is directly provided in IACT IRFs, in the case of LHAASO we convolved the expected background flux with the effective area calculated for hadrons, namely after the implementation of CR rejection cuts such that only residual background is accounted for. For the significance we here adopt the definition given by Li & Ma (1983), with and , being the ratio of the on-source to the off-source exposure; for our calculations, we assume .
@@
Concerning the condition on the background uncertainty, for all detectors we assume a 1% accuracy on the modeling (and hence subtraction) of the residual background and require a signal excess of at least five times this background systematic uncertainty, i.e. , following the approach adopted by ASTRI and CTA (Cherenkov Telescope Array Consortium et al., 2019).
Moreover, to account for the statistical fluctuations of the background, the number of background events is randomly extracted from a Poissonian distribution with mean value , and the result shown is averaged over 1000 realizations of the sensitivity estimation. Given the different observation strategy of the instruments, a benchmark observation time of 50 h is assumed for all IACTs while 1 year is considered for LHAASO, unless otherwise specified. Furthermore, IRFs at an average zenith angle of 20 are used for the calculation of the extended source sensitivities, given that public IRFs are available for all the observatories under investigation specifically at this value. The effects of IACT observations at different zenith angles are discussed in Appendix B. We consider here disk-like sources with eight different radial sizes, i.e. ; the choice of these specific sizes is mostly driven by an analog computation in our previous publication, see (Ambrogi et al., 2018). The instrument PSF affects the effective radius in which the signal is spread (), which we calculated as the convolution between the image of the source, described as a disk of radius , and the PSF, assumed to be a Gaussian function with its standard deviation. Note that depends on energy (see Fig. 1(a)), and consequently also .
For IACTs, the IRFs further depend on the offset from the center of the camera, as described in Appendix A.
In order to account for the variation of performance across the FoV, we construct for each IRF a data cube, namely defining the corresponding value at each energy (E) and position (x,y) in the FoV. The source flux is analogously binned, with an energy dependence given by the spectral assumption. We proceed then with the determination of by convolving the image with the PSF, considering the proper kernel for each point of the grid. We then compute the number of signal and background counts expected within by considering a weighted average of the effective area inside the on-region. With respect to the approach previously implemented in Ambrogi et al. (2018), this method allows us to account for the degradation of IRFs across the FoV and also to extend the computation to off-axis sources, i.e. centered at given distance from the center of the FoV. We discuss this point further in Appendix A, where we also present differential sensitivities at different offsets in Fig. 13.
4.1 Comparisons of sensitivities among different instruments
@@
On the other hand, the degradation due to the offset is restrained to a factor of a few units for all IACTs, as shown in Fig. 13. A comparative evaluation of the impact of these two effects (source extension and offset) is provided in Fig. 3 , where the sensitivities computed for different extensions () and at different offsets (), for IACTs, are compared to the sensitivity towards observations of a point source at the center of the camera (). Interestingly, ASTRI sensitivity is almost unaffected by the offset, mainly because of the small variation of its IRFs across the FoV, as it can be seen in Fig. 10. This guarantees the possibility of studying multiple sources in the same pointing without losing much in performance.
To test our methods, we compared our calculations with the sensitivity for point-like sources calculated by the CTA and ASTRI collaboration, finding consistent results (see Appendix C and Fig. 15). A consistency check among the resulting LHAASO-KM2A sensitivity and the sources listed in its first catalog can be performed, shown in Fig. 5(a) and further discussed later on.
@@ 5 Science cases
IACT and EAS measurements are complementary in many aspects: the former allow us to carry out precise spectroscopic and morphological studies of sources thanks to their superior energy and angular resolutions, while the large duty cycle and field of view of EAS arrays enable the deepest exposure to sources. The combination of these features will permit to investigate with an unprecedented level of accuracy some of the most relevant key science cases for the community. Specifically, the following developments are expected:
i) The improved angular resolution of IACTs will enable us to precisely locate the regions where particle interactions take place, within or around the accelerator, thus revealing information about the nature of the interaction itself (e.g. the presence of correlated emission with dense gas regions hints towards hadronic collision origin). This will further allow us to understand sources remaining unidentified in current VHE and UHE catalogs, as the first LHAASO catalog (Cao et al., 2023) or in the H.E.S.S. Galactic Plane Survey (HGPS), both of which are discussed in Sec. 5.1.
@@
iii) The large FoV of the next-generation IACTs will further allow us to better constrain the background around sources, facilitating the detection and the characterization of low-surface brightness sources such as molecular clouds, discussed in Secs. 5.3. Together with the improved angular resolution, this will permit to better model particle propagation around their sources: e.g., we expect the source size to change in energy in a leptonic scenario because of energy losses, as e.g. observed in the case of gamma-ray observations of the microquasar SS443 (MAGIC Collaboration et al., 2018; Abeysekara et al., 2018), while in hadronic scenarios the most energetic particles will travel loss-free to a further distance before interacting. An application of the propagation and escape concept certainly concerns Pulsar Wind Nebulae (PWNe) and their surrounding TeV halos (Giacinti et al., 2020), as we discuss in Sec. 5.4.
5.1 Catalogs follow-ups
Currently, both the HGPS (H. E. S. S. Collaboration et al., 2018) and the first LHAASO source catalog (Cao et al., 2023) contain more than half unidentified sources, requiring follow-up studies with improved-performance instruments. In these regards our calculations provide a clear assessment regarding the detection potential of each instrument and the time required to achieve a significant detection. We therefore proceed to the exploration of the two aforementioned source catalogs, in order to establish the most promising targets for each observatory.
5.1.1 The first LHAASO source catalog
After the announcement of the first 12 sources (Cao et al., 2021), a new catalog was recently published by the LHAASO Collaboration (Cao et al., 2023), reporting on the observation of 90 sources of both Galactic and extra-galactic nature. Among these, 32 represent new discoveries and 43 emitters are detected in the UHE domain with a test statistics () above 20 at 100 TeV: these sources are labeled with the letter in Cao et al. (2023). The first LHAASO catalog was obtained by analysing the WCDA and KM2A data-sets separately, and only 54 sources are significantly observed by both instruments. The differential spectra measured by the two instruments are fitted independently, each with a power-law shape and the best-fit parameters are reported in the catalog. Although a tentative association is proposed in Cao et al. (2023), the large extensions (see Fig. 4) and superposition of sources, together with a poor understanding of the diffuse emission at such high energies, make the identification of the sources challenging. IACTs will at least partly compensate these issues, provided that a reasonable exposure will be guaranteed to detect source fluxes.
@@ 5.1.1 The first LHAASO source catalog
We show in Fig. 5 the differential fluxes of the catalogued LHAASO sources at 3 TeV and 50 TeV, provided respectively in the WCDA and KM2A catalogs, as a function of the measured source extension (Cao et al., 2023). We compared them to the sensitivity of future IACTs at the same energies and for different exposure times. The sources are further divided depending on their culmination zenith angles, where red and blue points in the figure represent sources that culminate at a zenith below 30 at the latitude of CTA-South and CTA-North (which is almost at the same latitude of ASTRI), respectively.
With regards to WCDA sources, we find that almost all of the (so-defined) Southern sources will be accessible to CTA-South with 50 hr exposure, in line with the expectations. In fact, because the Southern sources fall at a higher zenith angle at the LHAASO site, a selection bias towards the brighter sources might be in place. In other words, because the LHAASO exposure is suppressed for sources observed at high-zenith angles (Aharonian et al., 2021a), most likely only the brightest UHE sources in the Southern hemisphere are included so far in the LHAASO sample. Many of the Northern hemisphere sources will be seen by CTA-North and ASTRI with 50 hr exposure, and almost all with a longer exposure around 300 hr.
With regards to KM2A sources, it appears that the extension of the sources might prevent in certain cases a clear detection with the considered observational setup. In particular, while many sources will be observed by CTA-South with 50 hr exposure, some will require a longer observation time, possibly up to h. Alongside, many of the detected KM2A-UHE sources will be detectable, allowing for a better characterization of these accelerators. Deep (300 hr) observations with ASTRI and CTA-North will also achieve a good-enough sensitivity to allow the follow-up of a handful of these interesting sources already in the next few years.
@@
The HGPS (H. E. S. S. Collaboration et al., 2018) is a survey performed by H.E.S.S. of the Galactic Plane (, ) containing 78 sources, detected down to 1% of the Crab flux. More than 60% of the sources are unidentified and 80% are extended. Moreover, only for 15% of sources (12 objects) energy spectra are consistent with exponential cut-off power laws (a pure power law is fitted otherwise), revealing a maximum cut-off energy in the HGPS at 19.2 TeV. In these respects, it appears clear that observations with next-generation instruments, more sensitive than H.E.S.S. at energies TeV, are needed in order to characterize the actual maximum energy achieved by these sources. While we expect CTA to improve the observations of basically all H.E.S.S. detected sources around 10 TeV, it is worth exploring the capability of next-generation IACTs in following up HGPS sources at higher energies. Despite the sky coverage of the Northern observatories and H.E.S.S. is not identical, a few sources are visible from both hemispheres, although culminating at different zenith angles depending on the site. In Fig. 6, we show the HGPS source fluxes at 10 and 50 TeV as a function of the corresponding extension. Note that flux values at 50 TeV are obtained by extrapolating the given spectral slope in the assumption of a pure power-law or an exponential-cutoff power-law with the tabulated parameters for each source. We also show the sensitivity of ASTRI for different exposures at a fixed zenith angle of and those of the two CTA observatories with 300 hr observations in a range of zenith angles between and , represented by a band. Further details about the expected CTA sensitivities with varying zenith are provided in Appendix B, showing that above few tens of TeV an improvement of about a factor 2 can be achieved in high-zenith angle observations, similarly to what was demonstrated in the case of MAGIC (Sommers & Elbert, 1987; Konopelko et al., 1999; MAGIC Collaboration et al., 2020). We expect analogous considerations to hold for ASTRI, but lacking of the proper IRFs at different culmination angles, we limit the ASTRI discussion to low-zenith angles sensitivities. Moreover, we include a color scale indicating the zenith angle at culmination for each source at the latitude of ASTRI: in particular the scale takes a white color for high-zenith angle observations, meaning a less solid comparison with the sensitivity here shown. As a result of the comparison, we find that with a long enough exposure ASTRI will allow to investigate almost the entire HGPS sample at 10 TeV. Similarly, good prospects are also expected for 50 TeV observations, where more than half of the sources will be in the reach of both ASTRI and CTA.

### v2

@@ 3.2 CTA and ASTRI IRFs
Fig. 1 also shows the IACT IRFs for on-axis observations (while IRFs for off-axis observations at different offsets from the FoV centre are shown in Appendix B) for both ASTRI and the two CTA observatories. The latter can be retrieved in publicly available zenodo repositories (ASTRI-Project, 2022; Cherenkov Telescope Array Observatory & Cherenkov Telescope Array
Consortium, 2021): because these are optimised for point-like sources, tailored event selection will need to be assessed in the future specifically for extended sources, and might provide improvements with respect to those adopted here. Also
nominal pointing is assumed in the simulations, with all telescopes pointing to the same direction.
In IACT observations, hadron-initiated showers constitute a background for photon-induced cascades, hence they need to be rejected after the data acquisition (Maier & Knapp, 2007). A first gamma/hadron separation is conducted based on the projected shape of the Cherenkov-light cone on the camera (Hillas, 1985): the high transverse momentum of secondary hadrons, in fact, induces a highly scattered distribution of hadronic showers at fixed energy of the primary. Nonetheless, an irreducible background of gamma-ray like hadron will survive the separation cuts. Such a residual hadronic emission is estimated at the level of the analysis, for example, by evaluating the background in a source-free region or by fitting a background template to the data in the region of interest (Berge et al., 2007).
Recently, 3D likelihood statistical methods have been implemented also in VHE data analysis, showing improved results especially for low surface brightness sources or for objects in crowded regions (Mohrmann et al., 2019; Vovk et al., 2018). The simplified approach used here is rather based on classical photon counting analysis, as discussed later in Sec. 4. We note that this choice will result into a more conservative estimation of source detectability.
A direct comparison among LHAASO, ASTRI, and CTA IRFs is now possible given that the same energy binning was set, namely 0.2 in logarithmic scale. The excellent angular resolution of IACTs, due to their pointing technique and camera, is visibly superior to the one of LHAASO-KM2A: in fact, the width of Point Spread Function (PSF) amounts to for IACTs vs for LHAASO. With regards to the effective area, instead, for EAS arrays, it coincides with the geometrical area, namely the instrument surface exposed to the shower, since shower reconstruction relies on the number of particles collected on ground. In the case of LHAASO, a surface of 1 km2 is covered by the KM2A. On the other hand, the effective area of IACTs is rather a convolution between the geometrical area and the Cherenkov light pool illuminated by the gamma-ray induced shower. The former depends on the specific layout of the experiment, namely by the relative distance among telescopes. For ASTRI, the foreseen layout covers an area of about m2 (Vercellone et al., 2022a), while CTA-North of about m2 and CTA-South of about m2 (Cherenkov Telescope Array Consortium et al., 2019). On the other hand, the Cherenkov pool of a primary photon interacting at a height of about 10 km spreads over a surface on ground of m2 (Aharonian et al., 2013): this value hence represents the effective area of an individual imaging telescope, which has to be rescaled to the number of telescopes when providing a description of the array. The effective areas shown in Fig. 1 are obtained after the gamma/hadron separation cuts, while no selection relative to the reconstructed event direction was applied333https://doi.org/10.5281/zenodo.5499840.
@@ 4 Sensitivity studies towards extended sources
The calculation of differential sensitivities towards the observation of extended sources is here explained. This is motivated by the need to perform improved spectral studies of the newly detected UHE sources, as well as of the many unidentified sources which have emerged in VHE sky surveys. We highlight that, while the knowledge of the integral sensitivity is usually sufficient to evaluate the overall detection capabilities of an instrument, the differential approach enables a rigorous investigation of the performance towards spectral studies. In particular, the presence of cut-offs in the energy distribution of sources makes the differential approach the only reliable assessment of detection prospects beyond its pivot energy. On the other hand, the integral approach would tend to overestimate them, particularly in the case that a cut-off has not yet been clearly assessed. We hence proceed with a detailed discussion about differential sensitivity studies, while we refer the interested reader to Appendix A for the integral sensitivity computation.
Following the approach described in Ambrogi et al. (2018), we define the minimum detectable flux by imposing that the following conditions are satisfied in each energy bin of the computation (Bernlöhr et al., 2013), namely:
@@
The number of signal events expected per energy bin is obtained by folding the gamma-ray effective area with the Crab nebula spectrum444This is described in the TeV regime as a power-law function, , where (TeV cm2 s)-1 is the flux normalisation at 1 TeV and (Aharonian et al., 2004).. However, it should be noted that the final sensitivity computation is independent of the assumed source flux. The number of background events is computed from the background rate, assumed to be constant in time. While the background rate is directly provided in IACT IRFs, in the case of LHAASO we convolved the expected background flux with the effective area calculated for hadrons, as explained in the previous section. For the significance, , we here adopt the definition given by Li & Ma (1983), with and , being the ratio of the on-source to the off-source exposure; for our calculations, we assume .
@@
Concerning the condition on the background uncertainty, for all detectors we assume a 1% accuracy on the modelling (and hence subtraction) of the residual background and require a signal excess of at least five times this background systematic uncertainty, namely , following the approach adopted by ASTRI and CTA (Cherenkov Telescope Array Consortium et al., 2019).
Moreover, to account for the statistical fluctuations of the background, the number of background events is randomly extracted from a Poissonian distribution with mean value , and the result shown is averaged over 1000 realisations of the sensitivity estimation. Given the different observation strategy of the instruments, a benchmark observation time of 50 h is assumed for all IACTs while 1 year is considered for LHAASO, unless otherwise specified. Furthermore, IRFs at an average zenith angle of 20∘ are used for the calculation of the extended source sensitivities, given that public IRFs are available for all the observatories under investigation specifically at this value. The effects of IACT observations at different zenith angles are discussed in Appendix C. We consider here disk-like sources with eight different radial sizes, namely ; the choice of these specific sizes is mostly driven by an analogue computation in our previous publication, see (Ambrogi et al., 2018). However, we limit the discussion to sources with radius smaller than , which is much smaller than the FoV of any considered-instrument. Otherwise, for source sizes comparable to the instrument FoV, the background estimation should account for additional systematic effects beyond the statistical noise considered in the presented calculations. The instrument PSF affects the effective radius in which the signal is spread (), which we calculated as the convolution between the image of the source, described as a disk of radius , and the PSF, assumed to be a Gaussian function with its standard deviation. Note that depends on energy (see Fig. 1(a)), and consequently also .
For IACTs, the IRFs further depend on the offset from the centre of the camera, as described in Appendix B.
In order to account for the variation of performance across the FoV, we construct for each IRF a data cube, namely defining the corresponding value at each energy (E) and position (x,y) in the FoV. The source flux is analogously binned, with an energy dependence given by the spectral assumption. We proceed then with the determination of by convolving the image with the PSF, considering the proper kernel for each point of the grid. We then compute the number of signal and background counts expected within by considering a weighted average of the effective area inside the on-region. With respect to the approach previously implemented in Ambrogi et al. (2018), this method allows us to account for the degradation of IRFs across the FoV and also to extend the computation to off-axis sources, that is centred at given distance from the centre of the FoV. We discuss this point further in Appendix B, where we also present differential sensitivities at different offsets in Fig. 16.
@@
On the other hand, the degradation due to the offset is restrained to a factor of a few units for all IACTs, as shown in Fig. 16. A comparative evaluation of the impact of these two effects (source extension and offset) is provided in Fig. 3 , where the sensitivities computed for different extensions () and at different offsets (), for IACTs, are compared to the sensitivity towards observations of a point source at the centre of the camera (). With regards to the sensitivity worsening induced purely by source extension (on-axis), we note that IACTs are more affected than EAS because of the intrinsic sharper PSF of the former. On the other hand, with regards to the offset, it is interesting to note that ASTRI sensitivity is almost unaffected by the offset, mainly because of the small variation of its IRFs across the FoV, as it can be seen in Fig. 13. This guarantees the possibility of studying multiple sources in the same pointing without losing much in performance.
To test our methods, we compared our calculations with the sensitivity for point-like sources calculated by the CTA and ASTRI collaboration, finding consistent results (see Appendix D and Fig. 18). A consistency check among the resulting LHAASO-KM2A sensitivity and the sources listed in its first catalogue can be performed, shown in Fig and 5(a) and further discussed later on.
@@ 5 Science cases
IACT and EAS measurements are complementary in many aspects: the former allow us to carry out precise spectroscopic and morphological studies of sources thanks to their superior energy and angular resolutions, while the large duty cycle and field of view of EAS arrays enable the deepest exposure to sources. The combination of these features will permit some of the most relevant key science cases for the community to be investigated with
an unprecedented level of accuracy. Specifically, the following developments are expected:
i) The improved angular resolution of IACTs will enable us to precisely locate the regions where particle interactions take place, within or around the accelerator, thus revealing information about the nature of the interaction itself (e.g. the presence of correlated emission with dense gas regions hints towards hadronic collision origin). This will further allow us to understand sources remaining unidentified in current VHE and UHE catalogues, as the first LHAASO catalogue (Cao et al., 2024a) or in the H.E.S.S. Galactic Plane Survey (HGPS), both of which are discussed in Sec. 5.1.
@@
iii) The large FoV of the next-generation IACTs will further allow us to better constrain the background around sources, facilitating the detection and the characterisation of low-surface brightness sources such as molecular clouds, discussed in Secs. 5.3. Together with the improved angular resolution, this will permit particle propagation around their sources to be modelled better: for example, we expect the source size to change in energy in a leptonic scenario because of energy losses, as observed in gamma rays from the microquasar SS433, for example (MAGIC Collaboration et al., 2018; Abeysekara et al., 2018; H. E. S. S. Collaboration et al., 2024), while in hadronic scenarios the most energetic particles will travel loss-free to a further distance before interacting. An application of the propagation and escape concept certainly concerns Pulsar Wind Nebulae (PWNe) and their surrounding TeV halos (Giacinti et al., 2020, for example), as we discuss in Sec. 5.4.
As a general remark for the following discussion, we mention that we do not explicitly report on the coordinates of the sources under investigation, for which we invite the interested reader to look at the detection papers by the respective collaborations.
5.1 Catalogue follow-ups
Currently, both the HGPS (H. E. S. S. Collaboration et al., 2018) and the first LHAASO source catalogue (Cao et al., 2024a) contain more than half unidentified sources, requiring follow-up studies with improved-performance instruments. We therefore proceed to the exploration of such source catalogues, in order to establish the most promising targets for the observatories under consideration.
5.1.1 The first LHAASO source catalogue
After the announcement of the first 12 sources (Cao et al., 2021), a new catalogue was recently published by the LHAASO Collaboration (Cao et al., 2024a), reporting on the observation of 90 sources of both Galactic and extra-galactic nature. Among these, 32 represent new discoveries and 43 are significantly detected in the UHE domain555With a test statistics (TS100) above 20 at 100 TeV, these sources are labelled with the letter in Cao et al. (2024a).. The first LHAASO catalogue was obtained by analysing the WCDA and KM2A data-sets separately, and only 54 sources are significantly observed by both instruments. The differential spectra measured by the two instruments are fitted independently, each with a power-law shape and the best-fit parameters are reported in the catalogue. Although a tentative association is proposed in Cao et al. (2024a), the large extensions (see Fig. 4) and superposition of sources, together with a poor understanding of the diffuse emission at such high energies, make the identification of the sources challenging. IACTs will at least partly compensate these issues, provided that a reasonable exposure will be guaranteed to detect source fluxes.
@@ 5.1.1 The first LHAASO source catalogue
We show in Fig. 5 the differential fluxes of the catalogued LHAASO sources at 3 TeV and 50 TeV, provided respectively in the WCDA and KM2A catalogues, as a function of the measured source extension (Cao et al., 2024a). We compared them to the sensitivity of future IACTs at the same energies and for different exposure times. The sources are further divided depending on their culmination zenith angles (), being the geographical latitude of the observatory, and the declination of the source), where red and blue points in the figure represent sources that culminate at a zenith below 30∘ at the latitude of CTA-South and CTA-North (which is almost at the same latitude of ASTRI), respectively.
With regards to WCDA sources, we find that almost all of the (so-defined) southern sources will be accessible to CTA-South with 50 hr exposure, in line with the expectations. In fact, because the southern sources fall at a higher zenith angle at the LHAASO site, a selection bias towards the brighter sources might be in place. In other words, because the LHAASO exposure is suppressed for sources observed at high-zenith angles (Aharonian et al., 2021b), most likely only the brightest UHE sources in the southern hemisphere are included so far in the LHAASO sample. Many of the northern hemisphere sources will be seen by CTA-North and ASTRI with 50 hr exposure, and almost all with a longer exposure around 300 hr.
With regards to KM2A sources, it appears that the extension of the sources might prevent in certain cases a clear detection with the considered observational setup. In particular, while many sources will be observed by CTA-South with 50 hr exposure, some will require a longer observation time, possibly up to h. Alongside, many of the detected KM2A-UHE sources will be detectable, allowing for a better characterisation of these accelerators. Deep (300 hr) observations with ASTRI and CTA-North will also achieve a good-enough sensitivity to allow the follow-up of a handful of these interesting sources already in the next few years.
@@
The HGPS (H. E. S. S. Collaboration et al., 2018) is a survey performed by H.E.S.S. of the Galactic Plane (, ) containing 78 sources, detected down to 1% of the Crab flux. More than 60% of the sources are unidentified and 80% are extended. Moreover, only for 15% of sources (12 objects) energy spectra are consistent with exponential cut-off power laws (a pure power law is fitted otherwise), revealing a maximum cut-off energy in the HGPS at 19.2 TeV. In these respects, it appears clear that observations with next-generation instruments, more sensitive than H.E.S.S. at energies TeV, are needed in order to characterise the actual maximum energy achieved by these sources. While we expect CTA to improve the observations of basically all H.E.S.S. detected sources around 10 TeV, it is worth exploring the capability of next-generation IACTs in following up HGPS sources at higher energies. Despite the sky coverage of the northern observatories and H.E.S.S. is not identical, a few sources are visible from both hemispheres, although culminating at different zenith angles depending on the site. In Fig. 6, we show the HGPS source fluxes at 10 and 50 TeV as a function of the corresponding extension. Note that flux values at 50 TeV are obtained by extrapolating the given spectral slope in the assumption of a pure power-law or an exponential-cutoff power-law with the tabulated parameters for each source. We also show the sensitivity of ASTRI for different exposures at a fixed zenith angle of and those of the two CTA observatories with 300 hr observations in a range of zenith angles between and , represented by a band. Moreover, we include a colour scale indicating the zenith angle at culmination for each source at the latitude of ASTRI: in particular the scale takes a white colour for high-zenith angle observations, meaning a less solid comparison with the sensitivity here shown. As a result of the comparison, we find that with a long enough exposure, ASTRI allows almost the entire HGPS sample to be investigated at 10 TeV. Similarly, good prospects are also expected for 50 TeV observations, where more than half of the sources will be in the reach of both ASTRI and CTA.
Further details about the expected CTA sensitivities with varying zenith are provided in Appendix C, showing that above few tens of TeV an improvement of about a factor 2 can be achieved in high-zenith angle observations, similarly to what was demonstrated in the case of MAGIC (Sommers & Elbert, 1987; Konopelko et al., 1999; MAGIC Collaboration et al., 2020). We expect analogous considerations to hold for ASTRI, but because of the lack of the official IRFs properly evaluated at different culmination angles, we limit the discussion concerning ASTRI to low-zenith angles sensitivities. We further note that the zenith angle was evaluated to impact the expected sensitivities of IACTs by an amount that depends on energy differently for CTA-S and CTA-N, as it can be seen in Fig. 17. In particular, the variation is quite pronounced for CTA-S at 10 TeV and CTA-N at 50 TeV, while it is more contained for CTA-N at 10 TeV.